Asset K has an expected return of 10 percent and a standard deviation of 28 percent. Asset L has an expected return of 7 percent and a standard deviation of 18 percent. The correlation between the assets is 0.40. What are the expected return and standard deviation of the minimum variance portfolio?
Answer #1

Answer:

First, we calculate the weight of each asset in minimum variance portfolio.

Weight of asset K = (18% 2 - 0.40 * 28% * 18%) / (28%2 + 18% 2 - 2 * 0.40 * 28% * 18%)

= 0.173666

Weight of asset L = 1 - 0.173666 = 0.826334

Hence:

Expected return of the portfolio = Expected return of K * Weight of K + Expected return of L * Weight of L

= 10% * 0.173666 + 7% * 0.826334

= 7.52%

Expected return of the portfolio = 7.52%

Standard deviation of two asset portfolio:

Standard deviation of portfolio = SQRT (0.173666 2 * 28% 2 + 0.826334 2 * 18% 2 + 2* 0.173666 * 0.826334 * 28% * 18% * 0.40)

= 17.40%

Standard deviation of portfolio = 17.40%